Maria is m years old. Sara is 5 years younger than Maria and 3 years older than Irina. What is the sum of the ages of all 3 girls?
A) 3m + 8
B) 3m + 2
C) 3m - 2
D) 3m - 8
E) 3m - 13
THanks!

Answer by edjones(8007)

sara=m-5
irina=m-8
m+m-5+m-8=3m-13 E
.
Ed